### v1

# Wav2Small: Distilling Wav2Vec2 to 72K parameters for Low-Resource Speech emotion recognition

Abstract
Speech Emotion Recognition (SER) needs high computational resources to overcome the challenge of substantial annotator disagreement. Today SER is shifting towards dimensional annotations of arousal, dominance, and valence (A/D/V). Universal metrics as the L2 distance prove unsuitable for evaluating A/D/V accuracy due to non converging consensus of annotator opinions. However, Concordance Correlation Coefficient (CCC) arose as an alternative metric for A/D/V where a model’s output is evaluated to match a whole dataset’s CCC rather than L2 distances of individual audios. Recent studies have shown that Wav2Vec2.0 / WavLM architectures outputing a float value for each A/D/V dimension achieve today’s State-of-the-art (SOTA) CCC on A/D/V. The Wav2Vec2.0 / WavLM family has a high computational footprint, but training tiny models using human annotations has been unsuccessful. In this paper we use a large Transformer SOTA A/D/V model as Teacher/Annotator to train 5 student models: 4 MobileNets and our proposed Wav2Small, using only the Teacher’s A/D/V predictions instead of human annotations. The Teacher model we construct also sets a new SOTA on the MSP Podcast dataset with valence CCC = $0.676$. We chose MobileNet-V4 / MobileNet-V3 as students, as MobileNet has been designed for fast execution times. We also propose Wav2Small – an architecture designed for minimal parameter number and RAM consumption. Wav2Small with an .onnx (quantized) of only $60KB$ is a potential solution for A/D/V on hearing aids, having only 72K parameters [^1] Convolution followed by BatchNorm then ReLU is a fused op in .onnx. vs 3.12M parameters for MobileNet-V4-Small.222© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
Index Terms—  speech emotion recognition, MobileNetV4
1 INTRODUCTION
Table 1: A/D/V model’s resources at 32bit non quantized .onnx. All onnx take 16KHz audio as input and include all parameters, preprocessing and LogMel Spectrogram extraction. Values from Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz.
5s audio input
Param.
Peak
RSS
MAC
Time CPU
Teacher $483.9$ $1929$ 149G $1089$
WavLM $318.6$ $1284$ 95G $688$
Dawn $165.3$ $697$ 55G $372$
MobileNetV4-L $31.87$ $257$ 2.2G $24$
MobileNetV4-M $10.38$ $94$ 1.7G $13$
MobileNetV3-S $3.14$ $22$ 400M $11$
MobileNetV4-S $3.12$ $36$ 417M $5$
Wav2Small $0.072$ $9$ 416M $9$
A need for SER in embedded devices of highly constrained resources, such as hearing aids, drives us to investigate small architectures for A/D/V. Neural Architecture Search has birthed useful architectures for categorical emotions SER [1]. However, SOTA A/D/V is dominated by Transformer models, such as Wav2Vec2 [2] or WavLM [3], both having a VGG7 input feature extractor. This VGG architecture is chosen because the skip connection of ResNet has high RAM footprint. In this paper we apply distillation to train 5 small student architectures for A/D/V. We evaluate three MobileNetV4 [4] students: (-Large, -Medium, -Small [^3] We use the Top-1 pretrained models from https://huggingface.co/timm/
mobilenetv4_conv_aa_large.e230_r448_in12k_ft_in1k
mobilenetv4_hybrid_medium.ix_e550_r384_in1k
mobilenetv4_conv_small.e1200_r224_in1k.
) as students for A/D/V. As well as a MobileNetV3-Small [5]444pretrained on Audioset., as well as our proposed architecture named Wav2Small.
1.1 Wav2Small
The paradigm of a VGG feature extractor followed by transformer layers has shown great performance for Speech Tasks [6]. Transformer pretrained architectures using time-samples input have been found superior than LogMel input for SER, yet LogMel input is still superior than time-samples for small architectures [7]. In this paper we propose Wav2Small: a VGG7 of only $13$ channels that uses vectorisation of time-frames (tokens) into the convolution-channel dimension to increase channels, shown in Listing 1.1. Wav2small does not downsample tokens, thus it has potential to become an inexpensive feature extractor for transformer architectures, of the Wav2vec2/WavLM family. Wav2small has equal Multiply Accumulations (MAC) with MobileNetV4-S albeit only with 72K parameters and less RAM, as shown in Table 1.
{listing}
[t!] {Verbatim}[commandchars=
{}] from torch import nn from torchlibrosa import Spectrogram, LogmelFilterBank
def conv(in=13): return nn.Sequential(nn.Conv2d(in,13,k=3,pad=1, bias=False), nn.BatchNorm2d(13), nn.ReLU())
class Wav2Small(): def ˙˙init˙˙(): self.vgg7 = nn.Sequential( Spectrogram(fft=64, hop=32), LogmelFilterBank(fft=64, mels=26), conv(in=1), conv(), conv(), nn.MaxPool2d(3, stride=2, pad=1), conv(), conv(), conv(), conv(), nn.Conv2d(13, 13, k=1)) self.lin = nn.Linear(169, 169) self.sof = nn.Linear(169, 169) self.adv = nn.Linear(169, 3)
def forward(x): ”’x: (batch, time˙samples)”’ x -= x.mean(1) variance = (x * x).mean(1) + 1e-7 x /= variance.sqrt() x = self.vgg7(x) batch, channel, token, mel = x.shape # channel is non contiguous to mel thus # reshapes tokens together as attention. x = x.reshape(batch, token, channel * mel) x = self.sof(x).softmax(1) * self.lin(x) x = x.sum(1) # Bayesian pool return self.adv(x) # arousal, domin., val. Wav2Small in PyTorch. Note that x.reshape vectorises neighbor tokens to a new token before Bayesian pool.
2 A / D / V Distillation
Inspired by two publicly available SOTA A/D/V models: The 12x Transformer Layer Wav2Vec2.0 [2] named Dawn and the 12x Transformer layer WavLM [3] that won the 2024 MSP Podcast Challenge [^5] https://lab-msp.com/MSP-Podcast_Competition/leaderboard.php for A/D/V. We propose to use them as annotators of A/D/V to train small student models.
2.1 Teacher
We notice that running the WavLM and also the Dawn baselines on the same audio and averaging their outputs A/D/V we obtain a Teacher model that outperforms both WavLM & Dawn in terms of CCC. We open source the weights of Teacher here [^6] https://huggingface.co/dkounadis/wav2small. To the best of our knowledge this Teacher defines a new SOTA on MSP Podcast valence with a $ccc=0.676$, as shown in Fig. 2. We use Teacher frozen, as teacher for distillation.
2.2 Data / Augmentation
Distillation for categorical emotions has been used for cross-lingual transfer [8] or as a pretrain / finetune paradigm [9]. Here we opt for simultaneous distillation on MSP Podcast and unlabeled / non-speech data. We use strong, always-on, mixup [10] thus clean MSP Podcast speech is never shown to the models. This enables long trainings without overfitting [11]. Extensive work on the most useful labels for SER [12] has revealed that soft labels as produced by a teacher-model provide better of human annotations [13], hence we use only the Teacher output A/D/V as ground truth and discard the MSP Podcast human labels.
The distillation of a student runs for 50M steps (around 2 weeks) with batch $=16$, SGD of fixed learning rate (no schedule) $=5e-5$, weight decay $=0$, momentum $=0$, on Nvidida RTX A4000 GPU, on PyTorch v2.4.0. The training audio is generated by mixing (on the fly during training): $0.64\times\text{audio}_{1}+0.41\times\text{audio}_{2}$ where audio1 is a random $0.4s$ up to $12s$ excerpt from a audiotrack of 12 action movies (in house dataset contains full movies with music & effects) and audio2 is drawn from one of three buckets: (either an Audioset, or MSP Podcast v1.7, or Cochlscene / environmental background of city, street, bus, etc. ambient noise. The training audio is then cyclically rotated (np.roll) as well as drop of random frequency-bands and re-normalising of lost energy to the remaining non-zero frequencies (metallic speech effect) and shouting [14]. This is the final signal given to Teacher and student. It can overflow the .wav $[-1,1]$ range due to shout.
Fig. 1: Arousal-Valence predictions for Crema-d test set, coloring according to Crema-d ground truth categories.
Fig. 2: CCC of Teacher, WavLM and Dawn (baselines), as well as the 5 students, on MSP Podcast v1.7 Test1.
2.3 CCC objective with quadrant correction
A/D/V datasets are annotated using 7-point Likert scales [15]. The odd number of Likert bins encourages the annotator to pick a side (positive / negative or neutral A/D/V). To leverage this information we introduce an auxiliary loss function: If Teacher A/D/V output is on different quadrant (in the A/D/V 3D space) from studnet’s A/D/V output, we penalise the student (L1 loss) to correct its prediction. This auxiliary loss vanishes if teacher and student A/D/V quadrants agree. In this case only the always-on CCC loss function [^7] https://github.com/bagustris/ccc_mse_ser [16] applies.
Fig. 3: Wav2Small discrepancy from Teacher for Japanese audio track of Harry Potter vol1. Not included for train.
Fig. 4: CCC of all tested models on IEMOCAP (audio of all 5 Sessions has been concatenated to a single dataset), not used for training.
3 Experiments
For this paper we train 5 students whose resources are shown in Table 1 with MobileNetV4-S having the fastest execution (5ms) and Wav2Small having the lowest parameters and RAM.
Authors in [2] trained a CNN14 [17] of $4.9M$ parameters using the human annotations of MSP Podcast v1.7. This resulted in very low CCC especially for valence, known as the most difficult dimension to recognize in SER: Their CNN14 achieved only $0.248$ valence CCC vs $0.638$ achieved by Dawn on MSP Podcast v1.7 test-1 split. Now Fig. 2 shows that all 5 students reach a valence CCC $>=0.37$. From the 5 students, MobileNets score slightly higher than (non pretrained) Wav2Small. Still, Wav2Small reaches significant CCC, e.g. arousal CCC$=0.66$ in MSP Podcast test-1, and $=0.56$ in IEMOCAP [18], used as an out of distribution dataset for evaluation. MobileNetV3-S shows lower CCCs of MobileNetV4-S and also has longer execution: $11s$ compared to $5s$ as show in Table 1 whose values are measured from .onnx running on Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz. Fig 3 shows that Wav2Small follows the Teacher arousal & valence faithfully on a language not seen in training.
All 5 students start with a LogMel layer (incorporated in .onnx via convolution). An extra conv is added to all MobileNets to upscale the 1-channel output of LogMel to 3-channel RGB, needed to feed the pretrained weights.
3.1 Frequency Bands of LogMel
We discovered that num fft bands $=64$ leads to the best CCC for MobileNetV4-M/S, MobileNetV3-S and Wav2Small. For MobileNetV4-L we use num FFT bands = $144$, because Num FFT$=64$ provides only $26$ mel bins, which causes a hihgly asymmetric aspect ratio (if we look at LogMel as an image) yielding empty-input for convolutions. The larger Num FFT reduces the time resolution and therefore MobileNetV4-L achieves only CCC $=0.59$ arousal in Fig. 2.
3.2 Vectorisation of Time as Channels
The power of Wav2Small comes from a large time resolution (tokens) and the vectorisation of neighboring tokens into the channel dimension creates $169$ ”channels” for the Bayesian pool. The joint processing of the vectorised tokens by the Bayesian pool, see Listing 1.1, boosts Wav2Small to achieve arousal CCC $=0.66$ in MSP Podcast, $=0.56$ in IEMOCAP: Fig 2/4 without pretraining. The valence dimension is the most difficult to predict: MobileNetV4-S obtains $0.42$ and Wav2Small $0.37$ yet in out-of-domain movie as in Fig. 3 Wav2Small does track the output of the Teacher faithfully.
Thinking of the lower valence achieved by all students compared to the Teacher, [19] has shown that Transformer models via ASR pretraining have learned linguistic cues to solve SER. Our 5 students do not posess linguistic cues, aiming towards an A/D/V model that is fair for many languages.
A model equipped with audio and text features, such as the model that won the 2024 SER classification challenge [20] is a future avenue, although under the risk of introducing language/accent unfairness from focusing at words instead of sound’s tonality.
In Fig 1 we see that A/D/V values cluster categorical emotions to some extent. One also notices that MobileNetV4-S & Wav2Small obtain an even larger dispersity along the valence axis. Therefore, an auxiliary teacher [^8] https://huggingface.co/3loi/SER-Odyssey-Baseline-WavLM-Categorical should be added to distillation for predicting also categorical emotions. Perhaps via CCC rather than cross-entropy. Observer variability and annotator agreement have been investigated across various annotation schemes [21] with findings revealing higher consensus when annotating A/D/V dimensions compared to annotating one out of six emotion categories. Thus we espy A/D/V to be the main avenue for SER [22] in the future.
4 CONCLUSION
We propose a 72K-parameter architecture named Wav2Small that reaches significant A/D/V score on both MSP Podcast and IEMOCAP datasets, while using only $9$ MB RAM compared to $36$ MB RAM used by MobileNetV4-S, while having equal MACs. Mobilenet-V4 and MobileNet-V3 reduce the time-sample resolution to a single: 960-dimensional-token / 1s audio, whereas Wav2Small provides 2450: 169-dimensional-token / 1s audio, as its output (before pooling). Hence Wav2Small is a potential replacement of the expensive input-feature-extractor of Transformer architectures, such as Wav2Vec2.0 and WavLM. We also propose a Teacher model that sets a new SOTA on MSP Podcast Test-1 achieving valence CCC $=0.676$.